# Ring Deconvolution Microscopy: An Exact Solution for Spatially-Varying Aberration Correction

### v3

@@ 2.7 Extensions and variations
Several extensions and variations of RDM, discussed below, may find use even when a system is not space-varying, when calibration data is not available, or if the system is spatially-varying, but not LRI.
@@
If aberrations are not space-variant, traditional deconvolution should perform as well as RDM and be more computationally efficient. However, the RDM pipeline can still provide value by providing a means for estimating the spherical aberration coefficient from the calibration image via Seidel fitting. With this coefficient, we can generate a synthetic center PSF and perform deconvolution. We call this procedure Seidel deconvolution, and find that it essentially denoises the PSF measurement since it finds the closest synthetic PSF to the measured one. In the results shown in Fig. 6, Seidel deconvolution resolves smaller features and gives a cleaner reconstructed image than traditional deconvolution.
Fitting synthetic PSFs for the purpose of deconvolution is not new; existing work has fit a variety of parametric models to the experimentally measured PSF including a 2D Gaussian distribution [40], Gaussian mixture model [41], Zernike basis [32, 33, 42, 43, 44, 45], spherical aberration diffraction model [46], and Seidel ray model [47, 48], but to our knowledge, the fitting of Seidel polynomials in the pupil function is novel.
@@
One intriguing extension of RDM is to perform deblurring without calibration data (blind deblurring). Such a situation may arise because measuring PSFs is difficult, inconvenient, or not available for images captured in the past. Blind deblurring involves joint estimation of the PSF and the deblurred object from a single image (see Methods 4.4).
In principal, blind deblurring can be applied to any spatially-varying system; however, it is computationally intensive, so we demonstrate here only the simple case of a shift-invariant system. Thus, only the spherical Seidel coefficient needs to be estimated. Results are shown in Fig. 6 along with non-blind deconvolution methods for comparison. The blind deconvolution result is similar to that of Seidel deconvolution, though cannot resolve the smallest features and has overly high contrast due to noise. This can be mitigated by regularization.
Future work extending this idea to spatially-varying systems could use DeepRD, which is more computationally efficient than ring deconvolution, to iteratively search over the space of deblurring networks and choose the network with the sharpest reconstruction.
Non-LRI systems
RDM is designed for shift-varying systems that are linear revolution-invariant (LRI), but even when the system is not rotationally symmetric, it may still be useful. In such cases, RDM will not provide a complete recovery of the underlying object. Instead, it can undo the components of the blur that are revolution-invariant, which follows from the linearity of the PSF model. This may be a useful trade-off of performance and computational burden, given that full shift-varying deblurring methods are often computationally intractable. It is also worth noting that the mathematical arguments made in Theorem 1 can be readily extended to other forms of symmetry, not just rotational symmetry. This could prove useful for reflection-mode microscopes, which often have a left-right symmetry due to reflection from a mirror, or even for light-sheet microscopes who’s field is symmetrically aberrated due to the light sheet’s symmetry about the beam waist.
@@ 3 Discussion
We develop a new pipeline for image deblurring called Ring Deconvolution Microscopy (RDM) which encompasses both an accurate and analytically-derived deblurring technique, ring deconvolution, and a fast alternative, Deep Ring Deconvolution (DeepRD). Like standard deconvolution microscopy, our methods only require a single calibration image; however, they offer space-varying aberration correction. We back RDM with a new theory of imaging under rotational symmetry, which we call linear revolution-invariance (LRI), and an accurate implementation of the LRI forward model (ring convolution). For RDM calibration we also develop a procedure for fitting Seidel aberration coefficients from a single calibration image of randomly-placed point sources. Finally, we verify the accuracy of our RDM deblurring methods in both simulation and experimentally over 3 diverse microscopy modalities.
Given our demonstration of its broad utility across different imaging modalities, we hope that RDM will ultimately replace deconvolution microscopy as standard practice in widespread applications from biology to astronomy. The notion of rotational symmetry is deeply embedded in the theory of optics, and as such, we find it a natural and acceptable modeling assumption that can enable significant improvements in basic image processing. We believe that RDM will find most use in systems that approach optical extremes such as miniature microscopes or ultra large FoV systems, but may also empower optical designers to simplify hardware knowing they have the ability to better correct for aberrations digitally.
@@ 4.1 Ring convolution and ring deconvolution
We begin with a primer on notation. Let $g(u,v)$ describe the object’s intensity at $(u,v)$ and $h(x,y;u,v)$ describe the space-varying PSF - the intensity at $(x,y)$ of the PSF generated by a point source at $(u,v)$. We further use the notation $\tilde{g}$ to denote the transformation of $g$ to polar coordinates. Then the final image intensity $f(x,y)$ of a linear optical system is formed by the superposition integral [49]:
@@
Traditional image deconvolution approximates the system as linear space-invariant (LSI), which means the PSF is the same at all positions in the FoV, $h(x,y;u,v)=h(x-u,y-v;0,0)$. This simplifies the forward model in Eq. (2) by reducing it to a convolution with a single PSF. This greatly simplifies the computation for forward and inverse problems, but at the cost of being inaccurate for space-varying aberrations.
@@
This ring-wise computation, wherein points at different radii are filtered heterogeneously, is consistent with the underlying intuition in LRI: the blur varies radially. Our first main result is an efficient, FFT-based inversion of the ring convolution.
@@
This problem can be solved efficiently via an iterative least squares solver using Algorithm 1 as a substep. A Fourier interpretation of ring convolution is provided in Appendix C. While the above results are all exact, the discrete time implementations of them have small, but nonzero errors due to discretization. For example, the polar transformation in Algorithm 1 requires a small amount of interpolation.
@@
Input: $N\times N$ pixel image $g$; PSFs along one radial line $h^{(j)}$, $j=1,\dots,K$; corresponding distances $r_{j}$, $j=1,\dots,K$ of each PSF from the center. Output: LRI blurred image $f$
@@ 4.2 Fitting Seidel Coefficients to PSFs
Ring (de)convolution algorithms require $h$, the collection of PSFs along one radial line in the FoV. Fortunately, there is a convenient and compact alternative to measuring these manually. The Seidel coefficients [34, 50] are a polynomial basis that can represent any rotationally symmetric system. We mathematically describe the form of these aberrations in Appendix B.
The estimation procedure for estimating the Seidel coefficients involves fitting them to a single, sparse image of a few randomly scattered point sources (e.g. fluorophores on a microscope slide)—such an image is usually easier to obtain than an image of an isolated point source in the center of the FoV. The presence of off-axis PSFs in the calibration image provides information about all of the aberration coefficients. Though it may be possible to fit these coefficients from a single off-axis PSF, we find that a few, randomly placed PSFs provides a robust fit. Let $r_{1},\ldots,r_{J}$ be the object-plane radii of the $J$ points in the calibration image. We then find the primary Seidel coefficients $\hat{\omega}$ whose generated PSFs best match the measured PSFs. Once again, this searching procedure is succinctly stated as an optimization problem,
@@
It has been shown that for LRI systems, the optimal fit $\hat{w}$ achieves a diminishing error [34]. Furthermore, the 5 primary Seidel coefficients index the dominant aberrations present in practical imaging systems: sphere, coma, astigmatism, field curvature, and distortion. For more complex aberrations it is possible to add higher-order Seidel coefficients to the fit. In practice, we fit these 5 coefficients via the ADAM optimizer [51] and obtain reasonable local minima even though the problem is nonconvex (see Fig. 5). Armed with the estimated Seidel coefficients, we can generate synthetic PSFs at any radius.
@@ 4.3 Deep Ring Deconvolution (DeepRD)
DeepRD is designed to incorporate the Seidel coefficients into the deblurring process in a parameter-efficient and interpretable manner. To that end, we propose a neural network architecture inspired by the physical LRI image formation model. The first key design element is to use a modified Hypernetwork [52], a network which predicts the weights of another, task-specific “primary” network. In DeepRD, an MLP-based hypernetwork takes in Seidel coefficients and produces a deep deblurring network that specifically works for the given coefficients. Our second key design element is the use of ringwise convolution kernels. Specifically, the hypernetwork produces CNN kernels for each radius which the primary network applies ring-by-ring. This replicates the revolution-invariance assumption central to ring deconvolution and eases the space-invariant constraint of typical convolutional kernels. Together, this design enables a neural network that is a fraction the size of a conventional U-Net with improved performance and generalization. To our knoweledge, we are the first to utilize learnable ring convolutions in the context of deep learning.
@@ 4.5 Experimental details
Experimental details for the micro-endoscopy experiment can be found in  [15].
4.5.1 Sample Preparation
Live tardigrades Tardigrades were mixed-staged adults of the eutardigrade species Hypsibius exemplaris Z151 (reclassified from Hypsibius dujardini in 2017), purchased from Sciento (Manchester, United Kingdom). Animals were cultured as described in [53]. A mixture of starved and nonstarved tardigrades were stained overnight with Invitrogen nucleic acid gel fluorescent stain, whose excitation and emission maxima are 502 nm and 530 nm, respectively. Individual stained tardigrades were then isolated onto a glass slide for imaging. Meanwhile, the non-fluorescent samples (USAF resolution targets, and rabbit liver tissue) were obtained imaging-ready on glass slides.
BPAE cells Bovine pulmonary artery endothelial cells were obtained from Thorlabs. They are labeled with with MitoTracker™ Red CMXRos and Alexa Fluor™ 488 Phalloidin.
4.5.2 Imaging
UCLA Miniscope. We used the UCLA Miniscope v3 with the Ximea MU9PM-MBRD 12 bit, 2.2 micron pixel sensor. Optically, the Miniscope is comprised of a gradient index objective and achromat tube lens; further details are provided in [1]. In order obtain the system PSFs, we imaged $1\mu$m fluorescent beads smeared on a glass slide. For deconvolution microscopy calibration, we repeatedly diluted the bead solution with isopropyl alcohol until we were able to sufficiently isolate a single bead, whereas for RDM calibration we used a single dilution and imaged a slide containing a sparse collection of beads. We used a custom 3D Prior stage controlled with Micromanager v1.4 and Pycromanager [54].
@@ 4.5.2 Imaging
4.5.3 Image processing
@@
All experimental images were captured and stored in a raw, unprocessed format (npy or tif). Miniscope images underwent hot pixel removal (detailed in the public code) and normalization prior to deblurring. These images were cropped afterward by 10 pixels in each dimension to remove edge artifacts. Multicolor images were downsampled by a factor of 2, separated into two channels and deblurred independently. After deblurring, the channels were recombined and globally contrast stretched for display. Pseudocoloring was done with ImageJ using the Green/Magenta LUT. These images were also cropped for edge artifacts. The details of the multimode fiber images can be found in [15]. The bead images were upsampled by 3x and convolved with a Gaussian kernel (3/2 pixel width) after deblurring. Neuron images were convolved with a Gaussian kernel (1 pixel width) after deblurring. This was done according to [15]. For simulation data, images were normalized before deblurring and cropped after deblurring. All displayed PSFs were globally contrast stretched for display.
4.5.4 Computation
@@
PSF generation for the simulation experiments was done by synthetically generating pupil functions with the given Seidel coefficients [50] (see Github). The remaining computation was done using Python on a single GPU, either a NVIDIA GeForce RTX 3090 or NVIDIA RTX A6000. For standard deconvolution the measured PSF was denoised through background subtraction and pixel-wise thresholding.
@@
All non-learning, iterative methods are solving linear least squares optimization problems (see (9)); we additionally add TV regularization to these and run them till convergence using an ADAM optimizer [51]. For each method, the hyperparameters—including learning rate and regularization strength—that provided the smallest loss and best qualitative results were used. For deconvolution we tried a variety of algorithms in addition to the iterative scheme, including Wiener filtering and Richardson-Lucy deconvolution, and used the best reconstruction, which was unsupervised wiener filtering [55].
@@
Open source implementations of ring convolution, polar transform, Seidel fitting, and ring deconvolution can be found in the codebase. Our intent is for this codebase to function as an easy-to-use library such that any practitioner with any imaging system can utilize RDM with little-to-no overhead.

### v5

@@ 2.5 Extensions and Variations
We have demonstrated our methods for rotational and lateral symmetry, but expect that the extension to other symmetries would be analogous. In addition, elements of RDM may find use even when the system is not space-varying, or when calibration data is not available.
@@
If aberrations are not space-variant, standard deconvolution should perform as well as RDM and be more computationally efficient. However, the first part of the RDM pipeline can still provide value by estimating the spherical aberration coefficient from a calibration image via Seidel fitting. With this coefficient, we can generate a synthetic center PSF and perform deconvolution. We call this procedure Seidel deconvolution, and find that it essentially denoises the PSF measurement since it finds the closest synthetic PSF to the measured one. In the results shown in Fig. 7, Seidel deconvolution resolves smaller features and gives a cleaner reconstructed image than standard deconvolution with the measured PSF. Fitting synthetic PSFs for the purpose of deconvolution is not new; existing work has fit a variety of parametric models to the experimentally measured PSF including a 2D Gaussian distribution [YangGaussian], Gaussian mixture model [samu], Zernike basis [Zheng:13, Shao:19, Maalouf:11, Bech, hanser2003phase, hanser04], spherical aberration diffraction model [Panka], and Seidel ray model [Wang_radial_ray, Simpkin], but to our knowledge, the fitting of Seidel polynomials in the pupil function is novel.
Figure 7: Variations of RDM for space-invariant systems. We compare results from our Seidel and blind deconvolution algorithms to that of standard deconvolution for images with a USAF test target and live tardigrades. Even for space-invariant systems, our methods outperform standard deconvolution by using a synthetic PSF, which prevents artifacts and loss of resolution from noise-based artifacts. Our blind deconvolution method usually estimates the spherical Seidel coefficient well; however, for certain images, the blind method can overestimate the coefficient, leading to oversharpening of the image, as in the USAF resolution chart.
@@
One intriguing extension of RDM is to perform deblurring without calibration data (blind deblurring). Such a situation may arise because measuring PSFs is difficult, inconvenient, or not available for images captured in the past. Blind deblurring involves joint estimation of the PSF and the deblurred object from a single image (see Methods 4.5). In principal, this can be done for any spatially-varying system; however, it is computationally intensive, so we demonstrate here only the simple case of a space-invariant system. Thus, only the spherical Seidel coefficient needs to be estimated. Results are shown in Fig. 7 along with non-blind deconvolution methods for comparison. The blind deconvolution result is similar to that of Seidel deconvolution, though cannot resolve the smallest features and has overly high contrast due to noise. Future work extending this idea to spatially-varying systems could use DeepRD, which is more computationally efficient than ring deconvolution, to iteratively search over the space of deblurring networks and choose the network with the sharpest reconstruction.
@@ 3 Discussion
In summary, we developed a new pipeline for image deblurring in rotationally-symmetric systems, called Ring Deconvolution Microscopy (RDM). RDM encompasses both an accurate and analytically-derived deblurring technique, and its fast alternative, Deep Ring Deconvolution (DeepRD). Like standard deconvolution microscopy, our methods only require a single calibration image; however, they offer space-varying aberration correction. We back RDM with a new theory of imaging under rotational symmetry, which we call linear revolution-invariance (LRI), and an implementation of the LRI forward model (ring convolution). For RDM calibration we also develop a procedure for fitting Seidel aberration coefficients from a single calibration image of randomly-placed point sources. In order to show generality of our ideas, we further derive, implement, and test an analogous method that exploits linear (instead of radial) symmetry, for applications in light-sheet microscopy. We verify the accuracy of our deblurring methods in both simulation and experimentally over four diverse microscopy modalities.
We hope that RDM will ultimately replace deconvolution microscopy as standard practice in widespread applications from biology to astronomy. We believe that RDM will find most use in systems that approach optical extremes such as miniature microscopes or large FoV systems, but may also empower optical designers to simplify hardware knowing they have the ability to better correct for aberrations digitally. RDM is well-suited to dynamic conditions in which the system or sample is changing, as long as the system calibration can be updated accordingly. For simpler cases that can be directly modeled, such as the time-varying deformation of an imaging fiber, it should be possible to update the initial calibrated PSFs with a theoretical model ( [ploschner2015seeing, wen2023single]).
We intend RDM to be a living, breathing tool with constant improvements to its speed and accuracy. Already we have seen a twenty-fold decrease in ring deconvolution’s runtime in preliminary experiments by parallelizing it over multiple GPUs using the novel Chromatix optical framework [chromatix_2023]. Moreover, the constant improvement in deep learning architecture, including better conditional models [chauhan2024brief], can also improve the performance of DeepRD in future. We plan to continually update our codebase with these improvements.
@@ 4.1 Ring Convolution and Ring Deconvolution
We begin with a primer on notation. Let $g(u,v)$ describe the object’s intensity at $(u,v)$ and $h(x,y;u,v)$ describe the space-varying PSF - the intensity at $(x,y)$ of the PSF generated by a point source at $(u,v)$. We further use the notation $\tilde{g}$ to denote the transformation of $g$ to polar coordinates. Then the final image intensity $f(x,y)$ of a linear optical system is formed by the superposition integral [goodman2005introduction]:
@@
standard image deconvolution approximates the system as linear space-invariant (LSI), which means the PSF is the same at all positions in the FoV, $h(x,y;u,v)=h(x-u,y-v;0,0)$. This simplifies the forward model in Eq. (2) by reducing it to a convolution with a single PSF. This greatly simplifies the computation for forward and inverse problems, but at the cost of being inaccurate for space-varying aberrations.
@@
where the $*_{\theta}$ operator indicates a one-dimensional convolution over the $\theta$ dimension. The $r$ arises in the deconvolution since we are integrating over object space $(u,v)$ in polar coordinates $(r,\theta)$. This ring-wise computation, wherein points at different radii are filtered heterogeneously, is consistent with the underlying intuition in LRI: the blur varies radially. Our first main result is an efficient, FFT-based inversion of the ring convolution.
@@
This problem can be solved efficiently via an iterative least squares solver using Algorithm 1 as a substep. A Fourier interpretation of ring convolution is provided in Appendix F. While the above results are all rigorous, the discrete time implementations of them have small, but nonzero errors due to discretization. For example, the polar transformation in Algorithm 1 requires a small amount of interpolation. As is the case with standard deconvolution, there are conditions for which ring convolution is not fully invertible and consequently ring deconvolution will not recover the sample exactly. The diffraction-limit, for example, manifests in PSFs whose frequency spectrum is bandlimited, rendering frequencies in the sample that are past the bandlimit irrecoverable. This can also happen if certain frequencies in the system’s transfer function are below the noise floor. In such cases ring deconvolution, because it is convex, will return an estimate of the sample that is closest in $l2$ norm to the true sample. Regularization can improve this result even further by leveraging prior knowledge about the sample in question.
@@
Input: $N\times N$ pixel image $g$; PSFs along one radial line $h^{(j)}$, $j=1,\dots,K$; corresponding distances $r_{j}$, $j=1,\dots,K$ of each PSF from the center.
Output: LRI blurred image $f$
@@ 4.2 Fitting Seidel Coefficients to PSFs
Ring (de)convolution algorithms require $h$, the collection of PSFs along one radial line in the FoV. Fortunately, there is a convenient and compact alternative to measuring these manually. The Seidel coefficients [born2013principles, voelz2011computational] are a polynomial basis that can represent any rotationally-symmetric system. We mathematically describe the form of these aberrations in Appendix D.
The estimation procedure for estimating the Seidel coefficients involves fitting them to a single, sparse image of a few randomly scattered point sources (e.g. fluorophores on a microscope slide). Such an image is usually easier to obtain than an image of an isolated point source in the center of the FoV. The presence of off-axis PSFs in the calibration image provides information about all aberration coefficients. Though it may be possible to fit these coefficients from a single off-axis PSF, we find that a few, randomly placed PSFs provides a robust fit. The locations of the point sources are not known a priori and are estimated via local peak detection. The optical center of the system is also needed in order to properly localize the PSFs; this can be found heuristically or by using common optical center finding algorithms such as [5206746].
Let $r_{1},\ldots,r_{J}$ be the object-plane radii of the $J$ points in the calibration image. We then find the primary Seidel coefficients $\hat{\omega}$ whose generated PSFs best match the measured PSFs. Once again, this searching procedure is succinctly stated as an optimization problem,
@@
It has been shown that for LRI systems, the optimal fit $\hat{w}$ achieves a diminishing error [born2013principles]. Furthermore, the 5 primary Seidel coefficients index the dominant aberrations present in practical imaging systems: sphere, coma, astigmatism, field curvature, and distortion. For more complex aberrations it is possible to add higher-order Seidel coefficients to the fit. See Appendix D.1 for an exploration of the higher-order terms. In practice, we fit these 5 coefficients via the ADAM optimizer [kingma2014adam] and obtain reasonable local minima even though the problem is nonconvex (see Fig. 5). Armed with the estimated Seidel coefficients, we can generate synthetic PSFs at any radius.
@@ 4.3 Deep Ring Deconvolution (DeepRD)
DeepRD is designed to incorporate the Seidel coefficients into the deblurring process in a parameter-efficient and interpretable manner. To that end, we propose a neural network architecture inspired by the physical LRI image formation model. The first key design element is to use a modified Hypernetwork [ha2017hypernetworks], a network which predicts the weights of another, task-specific “primary” network. In DeepRD, an MLP-based hypernetwork takes in Seidel coefficients and produces a deep deblurring network that specifically works for the given coefficients. Our second key design element is the use of ringwise convolution kernels. Specifically, the hypernetwork produces CNN kernels for each radius which the primary network applies ring-by-ring. This replicates the revolution-invariance assumption central to ring deconvolution and eases the space-invariant constraint of typical convolutional kernels. Together, this design enables a neural network that is a fraction the size of a conventional U-Net with improved performance and generalization. To our knoweledge, we are the first to utilize learnable ring convolutions in the context of deep learning.
@@
4.4 Sheet Convolution, Deconvolution, and the LSFM PSF Model
Here, we derive the efficient sheet convolution and deconvolution mathematically. The derivations are roughly similar to those for ring convolution, but with a different symmetry. Because of the similarities, our exposition here is more terse.
4.4.1 Sheet convolution and deconvolution
LSFM is a three-dimensional imaging modality and therefore the object and PSF will be a function of three space variables. Let $g(u,v,t)$ describe the object intensity at location $(u,v,t)$, and $h(x,y,z;u,v,t)$ describe the space-varying PSF due to Gaussian light-sheet illumination focused along the $u$ dimension (the choice of $u$ as the first spatial dimension is arbitrary). The superposition integral for a linear optical system can then be written as
[…]
analogously to (2). As before, we will incorporate the symmetry of LSFM to simplify the above display. The symmetry assumption states that the imaging optics are space-invariant, but the PSF varies in the $u$ direction due to the beam profile. Recall also that the total PSF is the product of the imaging PSF with the illumination PSF, which is varying. The following equation encodes these assumptions:
$h(x,y,z;u,v,t)=h(x,y-v,z-t;u,0,0).$ (12)
As before, plugging this into the linear forward model gives us sheet convolution:
[…]
where $*_{v,t}$ represents a 2D convolution along the $v$ and $t$ axes. From the equation we see that the image is the integral of two-dimensional convolutions of $y-z$ sheets of the object with $y-z$ sheets of the PSFs. To compute this integral, we only need to know the PSF at all values of $u$, and the $v$ and $t$ dimensions do not matter.
With a forward model in hand, sheet deconvolution can be solved exactly the same way as ring deconvolution: iterative least squares (Eq. (9)). The discretization and computational implementation of these algorithms mimic those of ring convolution and can be found in our codebase.
4.4.2 LSFM PSF model
The remaining component of LSFM deblurring strategy, is to obtain the necessary set of three-dimensional PSFs along one dimension. Similar to our earlier strategy for calibrating ring convolution/deconvolution, we can do so by imaging a single calibration volume containing randomly located, sparsely distributed point sources (e.g., fluorescent beads embedded in agarose). The resulting image stack will contain a random set of PSFs at different $u$ locations.
The simplest option is to directly estimate the PSFs at the missing $u$ locations by taking the convex combination of the two closest measured PSFs from the calibration stack. We use this strategy for the bead experiment in Fig. 6.
The second option is more akin to our Seidel fitting procedure; it involves parametrization of the spatially-varying PSFs with a unified, differentiable model. In this case we develop a modified version of popular Gibson-Lanni three-dimensional PSF model for LSFM [gibson1991experimental]. We will focus on our modifications of the model; further details about the Gibson-Lanni model and its variants are ubiquitous in the literature [gibson1991experimental, li2017fast]. Given a vector $p$ of system characteristics (e.g., sample refractive index) the Gibson-Lanni model calculates the optical path difference (OPD) between the ideal and experimental imaging systems. Integrating over this OPD gives the system’s three-dimensional PSF. Our LSFM PSF model takes this PSF and truncates its $z$ extent according to the light-sheet illumination thickness, thereby creating a spatially-varying PSF in $u$. Formally, let the Gibson-Lanni PSF be $h_{p}$, then our LSFM PSF
$h(x,y,z;u,v,t)=h_{p}(x-u,y-v,z-t)\frac{1}{\sigma(u)\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{z-t}{\sigma(u)}\right)^{2}},$ (14)
where
$\sigma(u)=\alpha\sqrt{1+(\beta u)^{2}}.$ (15)
The above equations arise from modeling the light-sheet as a Gaussian beam along $u$; its spread $\sigma(u)$ changes hyperbolically along the focus direction and its profile at a given $u$ is a Gaussian with variance $\sigma(u)^{2}$. We have two control knobs for it, $\alpha$, which controls the $z$ spread of the PSF at the thinnest section or waist, and $\beta$, which controls the rate that the spread increases as a function $u$. Given the calibration stack, we can optimize $\alpha$ and $\beta$ such that our PSF model produces PSFs close to the experimental ones. Since the PSF is a differentiable function of $(\alpha,\beta)$, they can be solved for using gradient-based iterative optimization, just like for Eq. (10). It is also possible to incorporate elements of $p$ into this optimization if they are not known a priori. This model is used to calibrate sheet deconvolution on the cell sample in Fig. 6.
In order to deploy the above models experimentally, one must ensure that the point sources are sufficiently sparse such that the PSFs are mostly non-overlapping. The exact point source locations are not important and can be estimated. Noisy PSFs are best handled with the Gibson-Lanni fitting method, which uses synthetic PSFs but still fits the measured PSFs well (see Appendix B). The interpolation method is more sensitive to noise but can be denoised effectively with thresholding and median filtering. The fitting procedures are detailed in our codebase, and follow the same general structure: first take a calibration stack of randomly-scattered point sources, estimate the locations of the PSFs using local peak finding algorithms, produce generated PSFs at those locations, and use the error between the generated and measured PSF to update the PSF model. In the case of interpolation, the last step is replaced by linearly interpolating the two closest measured PSFs to the desired location to create the generated PSF there.
@@ 4.6 Experimental Details
Experimental details for the micro-endoscopy experiment can be found in [Turcotte:20].
Sample Preparation
Live tardigrades Tardigrades were mixed-staged adults of the eutardigrade species Hypsibius exemplaris Z151 (reclassified from Hypsibius dujardini in 2017), purchased from Sciento (Manchester, United Kingdom). Animals were cultured as described in [lyons2024survival]. A mixture of starved and nonstarved tardigrades were stained overnight with Invitrogen nucleic acid gel fluorescent stain, whose excitation and emission maxima are 502 nm and 530 nm, respectively. Individual stained tardigrades were then isolated onto a glass slide for imaging. Meanwhile, the non-fluorescent samples (USAF resolution targets, and rabbit liver tissue) were obtained imaging-ready on glass slides.
BPAE cells Bovine pulmonary artery endothelial cells were obtained from Thermo Fisher. They are labeled with with MitoTracker™ Red CMXRos and Alexa Fluor™ 488 Phalloidin.
Beads 100nm Fluorescent beads were obtained from Polysciences (17150-10). For measuring the PSF in the light-sheet fluorescence microscope, we embedded the beads in 2% agarose with a final density ${}5x10^{-4}$ of the stock solution.
SU8686 cells SU8686 cells labeled with F-tractin-mRuby were embedded in soft bovine collagen, and then fixed before imaging with light-sheet fluorescence microscope.
Imaging
UCLA Miniscope. We used the UCLA Miniscope v3 with the Ximea MU9PM-MBRD 12 bit, 2.2 micron pixel sensor. Optically, the Miniscope is comprised of a gradient index objective and achromat tube lens; further details are provided in [aharoni]. In order obtain the system PSFs, we imaged $1\mu$m fluorescent beads smeared on a glass slide. For deconvolution microscopy calibration, we repeatedly diluted the bead solution with isopropyl alcohol until we were able to sufficiently isolate a single bead, whereas for RDM calibration we used a single dilution and imaged a slide containing a sparse collection of beads. We used a custom 3D Prior stage controlled with Micromanager v1.4 and Pycromanager [pinkard2021pycro].
@@
Light-sheet fluorescence microscope. We used a previously published setup for Field Synthesis [chang2019universal] that was operated without a ring mask, rendering it to a multi-directional selective plane illumination mSPIM [huisken2007even] system with a Gaussian sheet. The microscope equipped with 488 and 561 nm laser illumination, a Special Optics 28.5X/NA 0.67 illumination objective, and a Nikon 25X/NA 1.1 detection objective, and is controlled with a custom LabView program written by Coleman Technologies and is equipped with temperature control for long term live cell imaging.
Image processing
@@
All experimental images were captured and stored in a raw, unprocessed format (npy or tif). Miniscope images underwent hot pixel removal (detailed in the public code) and normalization prior to deblurring. These images were cropped afterward by 10 pixels in each dimension to remove edge artifacts. Multicolor images were downsampled by a factor of 2, separated into two channels and deblurred independently. After deblurring, the channels were recombined and globally contrast stretched for display. Pseudocoloring was done with ImageJ using the Green/Magenta LUT. These images were also cropped for edge artifacts. The details of the multimode fiber images can be found in [Turcotte:20]. The bead images were upsampled by 3x and convolved with a Gaussian kernel (3/2 pixel width) after deblurring. Neuron images were convolved with a Gaussian kernel (1 pixel width) after deblurring. This was done according to [Turcotte:20]. LSFM stacks were similarly cropped and contrast-stretched equally for each method for the purpose of display. For simulation data, images were normalized before deblurring and cropped after deblurring. All displayed PSFs were globally contrast stretched for display.
Computation
@@
PSF generation for the simulation experiments was done by synthetically generating pupil functions with the given Seidel coefficients [voelz2011computational] (see Github). Computation was done using Python on a single GPU, either a NVIDIA GeForce RTX 3090 or NVIDIA RTX A6000. For standard deconvolution the measured PSF was denoised through background subtraction and pixel-wise thresholding. For each $1024\times 1024$ image from the Miniscope and high NA multicolor systems, ring deconvolution took about 115 seconds and DeepRD took about $125$ milliseconds. For each $512\times 512$ image in simulation ring deconvolution took about 60 seconds and DeepRD took about $0.1$ seconds. For the $360\times 360$ images from the micro-endoscope, ring deconvolution took about 20 seconds. For a single $512\times 512\times 160$ volume from the LSFM system, sheet deconvolution took about 7 minutes.
@@
All non-learning, iterative methods are solving linear least squares optimization problems (see (9)); we additionally add TV regularization to these and run them till convergence using an ADAM optimizer [kingma2014adam]. For each method, the hyperparameters—including learning rate and regularization strength—that provided the smallest loss and best qualitative results were used. For deconvolution we tried a variety of algorithms in addition to the iterative scheme, including Wiener filtering and Richardson-Lucy deconvolution, and used the best reconstructions, which was either iterative deconvolution or unsupervised wiener filtering [Orieux:10].
@@
Open source implementations of ring convolution, polar transform, Seidel fitting, and ring deconvolution as well as the light-sheet extension methods can be found in the codebase. Our intent is for this codebase to function as an easy-to-use library such that any practitioner with any imaging system can utilize RDM with little-to-no overhead.